2x+9x-24/10+10x+6/6=26
Answer by ptaylor(2198)

(1) 2x+9x-24/10+10x+6/6=26
First, note that 6/6=1
And 2x+9x+10x=21x
Also 24/10=12/5
Now lets rewrite the equation:
(2) 21x-12/5+1=26 Multiply every term by 5 to get rid of the fraction
5(21x)-12(5/5)+1(5)=26(5) simplifying:
105x-12+5=130 add 7 to both sides:
105x-7+7=130+7
105x=137
x=137/105

ck
substitute in (2)
21x-12/5+1=26
21(137/105)-(12/5)+1=26
137/5-12/5+5/5=26
130/5=26
26=26